1. A bit string, 011111101111101111110, needs to be transmitted at the data link layer. What is the string actually transmitted after bit stuffing? Underline the additional bits you insert.
2. When should a selective repeat protocol be used versus a go-back-n protocol.
3. Specify the fields that make up a Point-to-Point protocol [PPP] frame.
4. Provides a simple parametric formula to quantify the size [in bytes] of a PPP frame.
5. A 100-km-long cable runs at the T1 data rate. The propagation speed in the cable is 2/3 the speed of light in vacuum. How many bits fit in the cable?
Answer #1

Q1

Bit Stuffing is used to distinguish data bits with the flag bits

Flag is 01111110 i.e(one 0 followed by 6 1's and last is 0)

so In bit stuffing if any data has continous 6 1's then we get confuse that either they are flag bits or data bits so we introduce a bit 0 after every 5 consecutive one and then transmit our final data

So given data is

011111101111101111110

Transmitted Data is ----> 011111010111110011111010

so we need to stuff 3 bits that are bolded

Q2

We use selective repeat over go-back n because of the following reasons

1. In Go-Back N we retransmits all the frame after the frame which gets lost whereas in selective repeat we only retransmits the lost frames

2. Due th above problem more bandwidth is wasted in retransmitting frames again and again in Go-Back N ARQ where as less bandwidth is needed in Selective Repeat

Because of these two reasons selective repeat is better

Q3

Fields in PPP frame are

  • Flag (1 byte) It is used to marks the beginning and the end of the frame.The bit pattern of the flag is 01111110.

  • Address (1 byte) which is set to provide the Ip address of the destination , If all are 1's that means it is abroadcast address

  • Control (1 byte ) is set to a constant value of 11000000.to show that frame doesn't contain any sequence number

  • Protocol − (1 or 2 bytes)---It define the type of data that will be present in the payload field.

  • Payload − This carries the data from the network layer. The maximum length of the payload field is 1500 bytes

  • FCS----->It is a 2 byte or 4 bytes frame check sequence for error detection. The standard code used is CRC (cyclic redundancy code)

These are filelds present in a PPP frame

Q5.

Propagation Delay = Length / speed

Propogation speed = 2/3 * 3 *10^8 m/sec = 2*10^8 m/sec = 2*10^5 km/sec

Propogation delay is 100/ 2*10^5= 0.5 *10 ^ - 3

Now

In One data rate data transfered = RTT *Data Rate

RTT= Round trip time --->2* Propogation Delay

RTT= 2*0.5 * 10^-3 = 1/1000 sec

So Total data carry by the cable = RTT * T1

= 1/1000 * T1 * 8 bits

here T1 the data rate and 1Byte = 8 bits

Thank You

If u like the answer please upvote it and if u have any doubt kindly comment on the answer before rating it .